OpenStudy (anonymous):

can someone check my result please? subtract 11x-5y from the sum of 7x+13y and -26x+19y. this is what I got. -30x+37y

OpenStudy (anonymous):

Can you show your work?

OpenStudy (anonymous):

(7x+13y)+(-26x+19y)= (7x-26x)+13y+19y)= -19x+32y and then I subtract 11x-5y (-19x+32y)-(11x-5y)= -19x+32y=11x+5y= -19x-11x+32y+5y= -30x+37y

OpenStudy (anonymous):

Looks good to me

OpenStudy (anonymous):

That's what I came up with I just wanted to see your work

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

anytime